OpenStudy (science0229):

Solve |x^2 - 4| = 2.

OpenStudy (anonymous):

x^2 - 4 = 2 x^2 = 6 x = sqrt6 x^2 - 4 = -2 x^2 = 2 x = sqrt2 Hence, x = sqrt2, sqrt6

OpenStudy (anonymous):

x+2/4=2

OpenStudy (science0229):

x^2-4=2 x^2=6 x=sqrt(6) or -sqrt(6) x^2-4=-2 x^2=2 x=sqrt(2) or -sqrt(2)